A wooden boat discovered just south of the Great Pyramid in Egypt as a carbon-14 / carbon -12 ratio that is 72.5% of that found in currently living organisms. How old is the boat?
Answer #1

carbon half life = 5730 yrs

decay constant = 0.693 / half life = 0.693 /5730 = 1.209 x 10^-4 yrs-1

A = 100 %

At = 72.5 %

t = 1/ K * ln (A/At)

t = 1/ (1.209 x 10^-4) * ln (100/72.5)

t = 2660 yrs

boat age = 2660 yrs